# Language-Based Bayesian Optimization Research Assistant (BORA)

### v1

@@ Abstract
Many important scientific problems involve multivariate optimization coupled with slow and laborious experimental measurements. These complex, high-dimensional searches can be defined by non-convex optimization landscapes that resemble needle-in-a-haystack surfaces, leading to entrapment in local minima. Contextualizing optimizers with human domain knowledge is a powerful approach to guide searches to localized fruitful regions. However, this approach is susceptible to human confirmation bias and it is also challenging for domain experts to keep track of the rapidly expanding scientific literature. Here, we propose the use of Large Language Models (LLMs) for contextualizing Bayesian optimization (BO) via a hybrid optimization framework that intelligently and economically blends stochastic inference with domain knowledge-based insights from the LLM, which is used to suggest new, better-performing areas of the search space for exploration. Our method fosters user engagement by offering real-time commentary on the optimization progress, explaining the reasoning behind the search strategies. We validate the effectiveness of our approach on synthetic benchmarks with up to 15 independent variables and demonstrate the ability of LLMs to reason in four real-world experimental tasks where context-aware suggestions boost optimization performance substantially.
@@ 1 Introduction
Figure 1: BORA framework. Icons from Flaticon (2025).
@@
Exploring large experimental design spaces requires intelligent navigation strategies because of the costly and time-consuming function evaluations involved. Bayesian optimization has been established as an optimal experimental design methodology in the scientific community across disciplines spanning chemistry Häse et al. (2021), solar energy production Andrés-Thió et al. (2024) and agronomy Zhang et al. (2024). BO can be used to efficiently navigate combinatorially large landscapes, and to identify promising solutions in an active-learning setting. Typically, BO uses a probabilistic surrogate model to approximate an expensive or unknown objective function $f$, i.e.,
$x^{*}=\underset{x\in\mathcal{X}}{argmax}\ f(x),$ (1)
with $f:\mathcal{X}\to\mathbb{R}$ defined on the compact subset $\mathcal{X}\subseteq\mathbb{R}^{d}$.
@@
This surrogate – often a Gaussian Process (GP) Garnett (2023) – subsequently undergoes Bayesian updates as new data about the design space are acquired, according to a predefined acquisition policy or function $\alpha(\cdot)$, allowing for the refinement of target predictions. In turn, this acquisition function suggests the next set of parameters for experiments with the highest expected utility, balancing exploration in new regions and exploitation in promising ones.
@@
Despite its successful application to a plethora of scientific tasks, BO is frequently characterized by slow initial optimization phases due to random or Latin Hypercube Poloczek et al. (2016) selections for initial samples. This can substantially slow the search for combinatorially large spaces. This highlights a fundamental challenge for standard BO; that is, the lack of inherent domain knowledge and contextual understanding of the problem at hand. Recently, BO variants have been proposed that are capable of injecting domain-specific knowledge into the search, either through structural problem characteristics Xie et al. (2023) or by using human expert knowledge Cissé et al. (2024). The latter approach, sometimes known as ‘Human-in-the-Loop (HIL)’, has drawn considerable recent attention, and it aims to infuse domain knowledge and human reasoning into BO workflows Adachi et al. (2024); Huang et al. (2022). By leveraging expert insights in the form of hypotheses Cissé et al. (2024), preferences Kristiadi et al. (2024) or specified priors over possible optima Hvarfner et al. (2022); Li et al. (2020), it is possible to enrich the optimization process and direct searches to fruitful regions much faster. Thus, HIL methods have shown increased effectiveness and efficiency compared with data-only approaches. In particular, hypothesis-based methods have shown gains in both performance and cost. Nonetheless, these HIL approaches can be human-capital resource-intensive because they require regular human interventions. Moreover, it is easy, even for domain experts, to lose track of the state-of-the-art in fast-moving research areas and ignore certain promising regions of the search space Ou et al. (2022).
@@
To address these challenges, we propose the use of Large language models (LLMs) Guo et al. (2024) as a facilitating framework in black-box optimization to enrich searches with domain knowledge. Specifically, we have coupled an LLM with standard BO in a hybrid optimization framework that automatically monitors and regulates the amount of domain knowledge needed when the search becomes ‘trapped’ in local minima. The algorithm capitalizes on the LLM’s inherent in-context learning (ICL) capacity and reasoning mechanism to suggest promising areas in the search space to sample from, in the form of hypotheses. LLMs have been employed recently to address limitations in core BO methodologies, as well as HIL variants Liu et al. (2024); Yin et al. (2024). LLMs have the capacity to encode vast amounts of domain-specific and general knowledge and have demonstrated the ability to reason about relatively complex tasks through in-context learning Kroeger et al. (2024); Xie et al. (2022) as well as in multidisciplinary domains such as chemistry Ramos et al. (2024). However, due to their numerically agnostic design, LLMs lag behind traditional BO methods in systematically balancing exploration versus exploitation, and have proved unreliable in many practical scenarios Huang et al. (2024). Recent attempts have been made to integrate LLMs with BO frameworks Mahammadli (2024); Yin et al. (2024) but thus far, these have been limited to small problem sizes, such as hyper-parameter optimization, or situations where the optimal solution is proximal to a special value Huang et al. (2024). LLM/BO hybrids could be prohibitively costly for more complex queries, particularly if the LLM is deployed for every iteration in the optimization.
@@
Here, we propose a language-based Bayesian Optimization Research Assistant, BORA, that enriches BO with domain knowledge and contextual understanding across a range of scientific tasks. We frame BORA as a hybrid optimization framework that boosts the performance of black-box optimizers by facilitating localizing areas of interest in the search space, informed by a knowledge-enriched LLM (Figure 1). A heuristic policy regulates the LLM involvement in the optimization process, thus optimally combining rigorous stochastic inference with LLM-generated insights within a feasible budget of LLM computation and API usage limits. During the intervention stage, the LLM uses its domain knowledge and reasoning capabilities to comment on the optimization progress thus far, highlighting patterns observed and forming hypotheses that may yield more rewarding solutions. It then tests these hypotheses by proposing new samples that maximize the target objective. BORA is also designed to provide an effective user-optimizer interaction through its dynamic commentary on the optimization process. This promotes deeper insights from the user and, in the future, the option to intervene; for example, by either reinforcing or overriding certain insights from the LLM. To our knowledge, this is the first time that a rigorous, dynamic synergy of black-box BO with LLMs has been proposed in this context. We evaluated BORA on various synthetic functions and a pétanque gaming model, as well as real scientific tasks in chemical materials design, solar energy production, and crop production. BORA demonstrated significant improvements in search exploration, convergence speed, and optimization awareness. Compared to earlier techniques, our method shows significant efficiency gains and generalization beyond hyper-parameter optimization, emphasizing its potential for tackling real-world tasks.
@@ 2 Related Works
To cope with non-convex optimization landscapes in science tasks, intelligent approaches have been proposed that focus on promising regions through adaptive exploration-exploitation strategies Shoyeb Raihan et al. (2024), or ‘smooth out’ the optimization landscape by enriching it with domain knowledge Ramachandran et al. (2020). Notable examples include local BO methods that restrict the search space, such as ZoMBI Siemenn et al. (2023), which aims to improve efficiency by focusing on local regions assumed to contain the optimum. Similarly, TuRBO Eriksson et al. (2019) uses multiple independent GP surrogate models within identified trust regions and a multi-armed bandit strategy Vermorel and Mohri (2005) to decide which local optimizations to continue. These approaches are well-suited to handling high-dimensional problems, but their potential is limited in small budgets and highly multimodal spaces due to a lack of built-in domain knowledge.
Incorporating domain knowledge into BO can significantly improve both its efficiency and its performance Adachi et al. (2024); Häse et al. (2021). DKIBO Xie et al. (2023) enhances BO’s acquisition function with structural knowledge from an additional deterministic surrogate model to enrich the GP’s approximation power. Others, such as ColaBO Hvarfner et al. (2024) and HypBO Cissé et al. (2024), allow users to inject their beliefs at the start to guide the optimization process. However, those methods keep the users’ beliefs static and cannot refine them as the optimization progresses, even if they are wrong. Meanwhile, other HIL methods rely on frequent user inputs Savage and del Rio Chanona (2023) and for robotic experiments Burger et al. (2020), for example, that run 24/7 in a closed-loop way, waiting for this human user input might become the rate-limiting step.
Recently, some studies have explored LLMs as standalone replacements for traditional optimizers due to their exceptional ability to solve complex problems in various domains Mirza et al. (2024); Nguyen and Grover (2024). Methods like LLAMBO Liu et al. (2024) and OPRO Yang et al. (2024) use the generative and ICL capabilities of LLMs to propose solutions to optimization problems directly. LLAMBO mimics BO’s structure and replaces its key components with LLMs. In OPRO, the LLM is iteratively prompted with the gathered optimization data as input and tasked to generate new solutions as output that are then evaluated. These methods are innovative but have focused so far on low-dimensional hyperparameter tuning and are not yet obviously suitable as a general framework for optimization tasks. Querying LLMs at all iterations also incurs a much larger computational and financial footprint than traditional BO algorithms. Moreover, standalone LLM optimizers depend on the static knowledge acquired during pre-training and lack the mathematical guarantees offered by traditional optimization algorithms.
@@
Our approach, BORA, shares similarities with the above studies by incorporating domain knowledge and adapting search mechanisms. However, BORA is distinguished by dynamically leveraging LLMs when they are most required, both for online hypothesis generation and for real-time commentary on optimization progress. Unlike static methods such as HypBO, which assume fixed human-injected soft constraints, our method refines the optimization trajectory based on the contextual insights given by the LLM. Moreover, BORA extends beyond previous hybrid approaches such as SLLMBO by introducing adaptive heuristics that intelligently modulate LLM involvement with BO to maximize optimization performance.
@@ 3 Methodology
The BORA optimization framework is illustrated in Figure 1. BORA is a fully automated hybrid BO-LLM synergy operating under a common GP whose parameters are updated accordingly as new points are sampled, either from BO or the LLM. A user-provided experiment description is used to initially contextualize the LLM which then warm-starts the optimization with proposed samples through its ICL capabilities. The optimization further progresses by alternating BO and LLM runs that are accordingly triggered by performance plateaus. Our proposed framework employs an adaptive heuristic policy to (a) assess the need to invoke the LLM, (b) determine the type of LLM intervention needed, and (c) update the frequency of those interventions as the optimization progresses. We use Reinforcement Learning (RL) terminology throughout the manuscript to describe our approach, however we use hand-crafted policy update rules, as learning generalized rules in the traditional sense Liu et al. (2022); Volpp et al. (2020) would be impractical in Bayesian scientific optimization settings Lee et al. (2020), which is the focus of this work. In its interventions, the LLM provides user interpretability through real-time commentary of the optimization progress and generates hypotheses for maximizing the target objective.
@@ 3.3 Actions
BORA leverages an adaptive heuristic policy detailed in Section 3.4 to choose one action from a set of three possible actions defined in the following paragraphs. The chosen action suggests at least one next point for evaluation, which is evaluated and added to the dataset. While the Vanilla BO action $a_{1}$ appends one sample $(x,y)$ to $\mathcal{D}_{t-1}$ at each step or iteration $t$, the LLM action $a_{2}$ (resp. $a_{3}$) adds $n_{\text{LLM}}\geq 1$ (resp. 2) samples, potentially adding multiple samples per step. Hereon, we distinguish between $t$, the step number, and $i$, the sample index at step $t$, and denote with $\mathcal{S}_{t}=\{x_{t}^{(i)}\}_{i=1}^{k}$ the set of $k$ points suggested by an action $a$ at step $t$.
@@ 3.3.2 $a_{2}$ LLM Comments and Suggests $n_{\text{LLM}}$ Points
A prompt containing the gathered data up to $\mathcal{D}_{t-1}$ and any previous comments $\mathcal{C}$ is given to the LLM. The LLM is then tasked to comment on the optimization progress in light of the new data and to update any previous hypotheses. The returned Comment contains the next points $\mathcal{S}_{t}=\{x_{t}^{(i)}\}_{i=1}^{n_{\text{LLM}}}\sim p(x|y>y_{\text{max}};\mathcal{D}_{t-1};\mathcal{C})$ to evaluate, which are then evaluated and added to the dataset to form $\mathcal{D}_{t}$.
3.3.3 $a_{3}$ LLM Comments and Selects 2 BO Points
Five candidate points $\{x_{\text{BO}}^{(j)}\}_{j=1}^{5}$ are generated by maximizing the acquisition function. A prompt containing $\mathcal{D}_{t-1}$, $\mathcal{C}$, and $\{x_{\text{BO}}^{(j)}\}_{j=1}^{5}$ are given to the LLM, which is then tasked to comment on the optimization progress in light of the new data and constrained to select the two most promising BO candidates that best align with its hypotheses for maximizing the target objective. The returned Comment holds the two selected points $\mathcal{S}_{t}=\bigl{\{}x_{t}^{(i)}\bigr{\}}_{i=1}^{2}\sim p(x\in\{x_{\text{BO}}^{(j)}\}_{j=1}^{5}|y>y_{\text{max}};\mathcal{D}_{t-1};\mathcal{C})$ that are then evaluated and added to the dataset.

### v2

# Abdoulatif Cissé1,2    Xenophon Evangelopoulos1,2    Vladimir V. Gusev3&Andrew I. Cooper1,2

@@ Abstract
Many important scientific problems involve multivariate optimization coupled with slow and laborious experimental measurements. These high-dimensional searches can be defined by complex, non-convex optimization landscapes that resemble needle-in-a-haystack surfaces, leading to entrapment in local minima. Contextualizing optimizers with human domain knowledge is a powerful approach to guide searches to localized fruitful regions. However, this approach is susceptible to human confirmation bias. It is also challenging for domain experts to keep track of the rapidly expanding scientific literature. Here, we propose the use of Large Language Models (LLMs) for contextualizing Bayesian optimization (BO) via a hybrid optimization framework that intelligently and economically blends stochastic inference with domain knowledge-based insights from the LLM, which is used to suggest new, better-performing areas of the search space for exploration. Our method fosters user engagement by offering real-time commentary on the optimization progress, explaining the reasoning behind the search strategies. We validate the effectiveness of our approach on synthetic benchmarks with up to 15 variables and demonstrate the ability of LLMs to reason in four real-world experimental tasks where context-aware suggestions boost optimization performance substantially.
@@ 1 Introduction
Figure 1: The BORA framework. Icons from Flaticon (2025).
@@
Exploring large experimental design spaces requires intelligent navigation strategies because of the costly and time-consuming function evaluations involved. Bayesian optimization has been established as an optimal experimental design methodology across disciplines spanning chemistry Häse et al. (2021), solar energy production Andrés-Thió et al. (2024) and agronomy Zhang et al. (2024). BO can be used to efficiently navigate combinatorially large landscapes, and to identify promising solutions in an active-learning setting. Typically, BO uses a probabilistic surrogate to approximate an expensive or unknown objective function $f$ while iteratively searching for a maximizer,
$x^{*}=\operatorname*{argmax}_{x\in\mathcal{X}}f(x),$ (1)
with $f:\mathcal{X}\to\mathbb{R}$ defined on the search domain $\mathcal{X}\subseteq\mathbb{R}^{d}$.
@@
This surrogate – often a Gaussian Process (GP) Garnett (2023) – subsequently undergoes Bayesian updates as new data about the design space are acquired, according to a predefined acquisition policy or function $\alpha(\cdot)$, allowing for the refinement of predictions of the objective $f(x)$. In turn, this acquisition function suggests the next set of parameters for experiments with the highest expected utility, balancing exploration in new regions and exploitation in promising ones.
@@
Despite its successful application to a plethora of scientific tasks, BO is frequently characterized by slow initial optimization phases due to random or Latin Hypercube Poloczek et al. (2016) selection of initial samples. This can slow the search for combinatorially large spaces substantially. This highlights a fundamental challenge for standard BO; that is, the lack of inherent domain knowledge and contextual understanding of the problem at hand. Recently, BO variants have been proposed that are capable of injecting domain-specific knowledge into the search, either through structural problem characteristics Xie et al. (2023) or by using human expert knowledge Cissé et al. (2024). The latter approach, sometimes known as ‘Human-in-the-Loop (HIL)’, has drawn considerable recent attention, and it aims to infuse domain knowledge and human reasoning into BO workflows Adachi et al. (2024); Huang et al. (2022). By leveraging expert insights in the form of hypotheses Cissé et al. (2024), preferences Kristiadi et al. (2024) or specified priors over possible optima Hvarfner et al. (2022); Li et al. (2020), it is possible to enrich the optimization process and to direct searches to fruitful regions much faster. Thus, HIL methods have shown increased effectiveness and efficiency compared with data-only approaches. In particular, hypothesis-based methods have shown gains in both performance and cost. Nonetheless, these HIL approaches can be human-capital resource-intensive because they require regular human interventions. Moreover, it is easy, even for domain experts, to lose track of the state of the art in fast-moving research areas and to ignore certain promising regions of the search space Ou et al. (2022).
@@
To address these challenges, we propose the use of Large Language Models (LLMs) Guo et al. (2024) as a facilitating framework in black-box optimization to enrich searches with domain knowledge. Specifically, we have coupled an LLM with standard BO in a hybrid optimization framework that automatically monitors and regulates the amount of domain knowledge needed when the search becomes ‘trapped’ in local minima. The algorithm capitalizes on the LLM’s inherent in-context learning (ICL) capacity and reasoning mechanism to suggest, in the form of hypotheses, promising areas in the search space from which to sample. LLMs have been employed recently to address limitations in core BO methodologies, as well as HIL variants Liu et al. (2024); Yin et al. (2024). LLMs have the capacity to encode vast amounts of domain-specific and general knowledge and have demonstrated the ability to reason about relatively complex tasks through in-context learning Kroeger et al. (2024); Xie et al. (2022) as well as in multidisciplinary domains such as chemistry Ramos et al. (2024). However, due to their numerically agnostic design, LLMs lag behind traditional BO methods in systematically balancing exploration versus exploitation, and have proved unreliable in many practical scenarios Huang et al. (2024). Recent attempts have been made to integrate LLMs with BO frameworks Mahammadli (2024); Yin et al. (2024) but thus far, these have been limited to small problem sizes, such as hyper-parameter optimization, or situations where the optimal solution is proximal to a special value Huang et al. (2024). Also, LLM/BO hybrids could be prohibitively costly for more complex queries, particularly if the LLM is deployed for every iteration in the optimization.
@@
Here, we propose a language-based Bayesian Optimization Research Assistant, BORA, that enriches BO with domain knowledge and contextual understanding across a range of scientific tasks. We frame BORA as a hybrid framework that augments surrogate-based optimizers with uncertainty estimates by localizing areas of interest in the search space, guided by a knowledge-enriched LLM (Figure 1). A heuristic policy regulates the LLM involvement in the optimization process, adaptively balancing rigorous stochastic inference with LLM-generated insights within a feasible budget of LLM computation and API usage limits. During the intervention stage, the LLM uses its domain knowledge and reasoning capabilities to comment on the optimization progress thus far, highlighting patterns observed and forming hypotheses that may yield more rewarding solutions. It then tests these hypotheses by proposing new samples that maximize the target objective. BORA is also designed to provide an effective user-optimizer interaction through its dynamic commentary on the optimization process. This promotes deeper insights from the user and, in the future, the option to intervene; for example, by either reinforcing or overriding certain insights from the LLM. To our knowledge, this is the first time that a rigorous, dynamic synergy of black-box BO with LLMs has been proposed in this context. We evaluated BORA on various synthetic functions and a pétanque gaming model, as well as real scientific tasks in chemical materials design, solar energy production, and crop production. BORA demonstrated significant improvements in search exploration, convergence speed, and optimization awareness. Compared to earlier techniques, our method shows significant efficiency gains and generalization beyond hyper-parameter optimization, emphasizing its potential for tackling real-world tasks.
@@ 2 Related Works
To cope with non-convex optimization landscapes in science tasks, intelligent approaches have been proposed that focus on promising regions through adaptive exploration-exploitation strategies Shoyeb Raihan et al. (2024), or ‘smooth out’ the optimization landscape by enriching it with domain knowledge Ramachandran et al. (2020). Notable examples include local BO methods that restrict the search space, such as ZoMBI Siemenn et al. (2023), which aims to improve efficiency by focusing on local regions assumed to contain the optimum. Similarly, TuRBO Eriksson et al. (2019) uses multiple independent GP surrogate models within identified trust regions and a multi-armed bandit strategy Vermorel and Mohri (2005) to decide which local optimizations to continue. These approaches are well-suited to handling high-dimensional problems, but their potential is perhaps more limited in small budgets and highly multimodal spaces due to a lack of built-in domain knowledge.
Incorporating domain knowledge into BO can improve both its efficiency and its performance Adachi et al. (2024); Häse et al. (2021). DKIBO Xie et al. (2023) enhances BO’s acquisition function with structural knowledge from an additional deterministic surrogate model to enrich the GP’s approximation power. Others, such as ColaBO Hvarfner et al. (2024) and HypBO Cissé et al. (2024), allow users to inject their beliefs at the start to guide the optimization process. However, those methods keep the users’ beliefs static and cannot refine them as the optimization progresses, even if they are wrong. Meanwhile, other HIL methods rely on frequent user inputs Savage and del Rio Chanona (2023) and for robotic experiments Burger et al. (2020), for example, that run 24/7 in a closed-loop way, waiting for this human user input might become the rate-limiting step.
Recently, some studies have explored LLMs as standalone replacements for traditional optimizers due to their exceptional ability to solve complex problems in various domains M. Bran et al. (2024); Nguyen and Grover (2024). Methods like LLAMBO Liu et al. (2024) and OPRO Yang et al. (2024) use the generative and ICL capabilities of LLMs to propose solutions to optimization problems directly. LLAMBO mimics BO’s structure and replaces its key components with LLMs. In OPRO, the LLM is iteratively prompted with the gathered optimization data as input and tasked to generate new solutions as output that are then evaluated. These methods are innovative but have focused so far on low-dimensional hyperparameter tuning and are not yet obviously suitable as a general framework for optimization tasks. Querying LLMs at all iterations also incurs a larger computational and financial footprint than traditional BO algorithms, particularly if reasoning models are used. Standalone LLM optimizers also lack the mathematical guarantees offered by traditional optimizers such as BO.
@@
Our approach, BORA, shares similarities with the above studies by incorporating domain knowledge and adapting search mechanisms. However, BORA is distinguished by leveraging LLMs when they are most required, for online hypothesis generation and for real-time commentary on optimization progress. Unlike static methods such as HypBO, which assume fixed human-injected soft constraints, our method refines the optimization trajectory based on the contextual insights given by the LLM. Moreover, BORA extends beyond previous hybrid approaches such as SLLMBO by introducing adaptive heuristics that intelligently modulate LLM involvement with BO to maximize optimization performance.
@@ 3 Methodology
The BORA optimization framework is illustrated in Figure 1. It is an automated hybrid BO-LLM synergy operating under a common GP whose parameters are updated as new points are sampled, either from BO or the LLM. A user-provided experiment description is used to initially contextualize the LLM which then warm-starts the optimization with proposed samples through its ICL capabilities. The optimization progresses by alternating BO and LLM runs that are accordingly triggered by performance plateaus. Our proposed framework employs an adaptive heuristic policy to (a) assess the need to invoke the LLM, (b) determine the type of LLM intervention needed, and (c) update the frequency of those interventions as the optimization progresses. We use Reinforcement Learning (RL) terminology in the manuscript to describe our approach, but we use hand-crafted policy update rules because learning generalized rules in the traditional sense Liu et al. (2022); Volpp et al. (2020) would be impractical in Bayesian scientific optimization settings Lee et al. (2020), which is the focus of this work. In its interventions, the LLM provides user interpretability via real-time commentary of the optimization progress and generates hypotheses to maximize the objective.
@@ 3.3 Actions
BORA leverages an adaptive heuristic policy detailed in Section 3.4 to choose one action from a set of three possible actions defined in the following paragraphs. The chosen action suggests at least one next point for evaluation, which is evaluated and added to the dataset. While the Vanilla BO action $a_{1}$ appends one sample $(x,y)$ to $\mathcal{D}_{t-1}$ at each step or iteration $t$, the LLM actions $a_{2}$ and $a_{3}$ add $n_{\text{LLM}}~\text{and}~n_{\text{LBO}}\geq 1$ samples, respectively. Hereon, we distinguish between $t$, the step number, and $i$, the sample index at step $t$, and denote with $\mathcal{S}_{t}=\{x_{t}^{(i)}\}_{i=1}^{k}$ the set of $k$ points suggested by an action $a$ at step $t$.
@@ 3.3.2 $a_{2}$ LLM Comments and Suggests $n_{\text{LLM}}$ Points
A prompt containing the gathered data up to $\mathcal{D}_{t-1}$ and any previous comments $\mathcal{C}$ is given to the LLM. The LLM is then tasked to comment on the optimization progress in light of the new data and to update any previous hypotheses. The returned Comment contains the next points $\mathcal{S}_{t}=\{x_{t}^{(i)}\}_{i=1}^{n_{\text{LLM}}}\sim p(x|y>y_{\text{max}};\mathcal{D}_{t-1};\mathcal{C})$ chosen by the LLM, which are then evaluated and added to the dataset to form $\mathcal{D}_{t}$.
3.3.3 $a_{3}$ LLM Comments and Selects $n_{\text{LBO}}$ BO Points
$a_{3}$ is a non-myopic, ICL step that focuses the LLM’s attention on a high-quality set of $n_{\text{BO}}$ candidate points $\{x_{\text{BO}}^{(j)}\}_{j=1}^{n_{\text{BO}}}$, which are generated by maximizing the acquisition function. A prompt containing $\mathcal{D}_{t-1}$, $\mathcal{C}$, and $\{x_{\text{BO}}^{(j)}\}_{j=1}^{n_{\text{BO}}}$ are given to the LLM, which is then tasked to comment on the optimization progress in light of the new data and constrained to select the $n_{\text{LBO}}$ most promising BO candidates that best align with its hypotheses for maximizing the target objective. Setting $n_{\text{BO}}=5$ and $n_{\text{LBO}}=2$ empirically showed to offer enough diversity of hypothesized optima locations and ensure competitive performance overall. The returned Comment holds the $n_{\text{LBO}}$ selected points $\mathcal{S}_{t}=\bigl{\{}x_{t}^{(i)}\bigr{\}}_{i=1}^{n_{\text{LBO}}}\sim p(x\in\{x_{\text{BO}}^{(j)}\}_{j=1}^{n_{\text{BO}}}|y>y_{\text{max}};\mathcal{D}_{t-1};\mathcal{C})$ that are then evaluated and added to the dataset.